OpenStudy (anonymous):

2(z + 1) - 4(z - 4) < (or equal to) 2(z + 5) solve for z?

OpenStudy (anonymous):

2z + 2 - 4z + 16 <= 2z + 10 -2z + 18 <= 2z + 10 +2z -10 +2z +10 8<= 4z /4 /4 2 <= z So the answer it 2

OpenStudy (anonymous):

Thank yOU!!!!!!!!!!!!!!!!!!!!! :D :D :D you are VERY MUCH appreciated ;)

OpenStudy (anonymous):

Hehe no problem :)

OpenStudy (anonymous):

Hey how come its plus sixteen?

OpenStudy (anonymous):

hold let me fix that, I made a mistake ;(

OpenStudy (anonymous):

okay...

OpenStudy (anonymous):

Ok nvm I did do it right, remember 2 negatives make a positive

OpenStudy (anonymous):

-4 x -4 = 16

OpenStudy (anonymous):

alright i get... thank you ... again :P

OpenStudy (anonymous):

No prob